OpenStudy (anonymous):

The area of a circle is 113.04 in. 2. What is the radius of the circle? 36 inches 18 inches 12 inches 6 inches

OpenStudy (calculusxy):

\[Area = \pi r^2\] Typically, people use the value of pi as 3.14 and we now the area is 113.04in^2. \[113.04inches^2 = 3.14r^2\]

OpenStudy (calculusxy):

Now you need to divide both sides by 3.14 to isolate r and to find the radius. \[\frac{ 113.04inches^2 }{ 3.14 } = \frac{ 3.14r^2 }{ 3.14 } => 36 = r^2\]

OpenStudy (anonymous):

a

OpenStudy (calculusxy):

Not really

OpenStudy (calculusxy):

Since r is squared, that means that we need to find the square root of 36.

OpenStudy (calculusxy):

\[\sqrt36 = r\]

OpenStudy (calculusxy):

Did you get your answer?

OpenStudy (anonymous):

d

OpenStudy (calculusxy):

Yes :)